Overall status: C config concludes on the 18th; weather issues toward the end of the year so a bit of a crunch to finish high priority commitments. Spate of issues (most recent the impressive sustained cold caused a spate of cryo leaks over the new year-limited high frequency observing over a period). Next week is the AAS - no meeting planned.
07Jan: Call for Proposals for February 2011 Call (first for new semester based system; tentative)
09-13 Jan: AAS with special NRAO sessions: Perley: 7 presentations scheduled for the EVLA special session.
Notes: 3c454.3 flux monitoring on-going; meeting this week(?); would like discussion of whether we are achieving theoretical noise at high frequencies as had been confirmed for low (recent issues seen at Ka band - Butler).
High Frequency: McMullin/Dhawan: Reference pointing with time averaging tests.
High Frequency: Claussen: Weak calibrator tests; needed.
Mid Frequency (Dhawan):
Low Frequency (Momjian):
Notes: Meeting this week.
Spectral Line (Goss/Ott):
Notes: Zeeman test follow-up.
Spectral Line: Momjian/Sowinski: Follow-up on Zeeman tests
Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
Notes: Discussions last week with Ransom; plans to piggy-back on phased-array testing in January; try to use CBE to create initial file (discussion on type of SDM, switching between de-dispersion mode and correlator mode). Paul to come for conversion from VDIFF to DSPSR format.